OpenStudy (anonymous):

Solve for x: 5x^2-10x+28=3+20x-4x^2 simplify, using radicals as needed. express complex numbers in terms of i.

OpenStudy (saifoo.khan):

Add like terms.

OpenStudy (saifoo.khan):

9 x^2-30 x+25 = 0

OpenStudy (saifoo.khan):

x = 5/3

jimthompson5910 (jim_thompson5910):

\[\large 5x^2-10x+28=3+20x-4x^2\] \[\large 5x^2-10x+28-3-20x+4x^2=0\] \[\large 9x^2-30x+25=0\] \[\large (3x-5)^2=0\] \[\large 3x-5=\sqrt{0}\] \[\large 3x-5=0\] \[\large 3x=5\] \[\large x=\frac{5}{3}\]